OpenStudy (anonymous):

Solve y = 3x – 12 by the substitution method. 2x – 3y = -13

OpenStudy (luigi0210):

sub in y: 2x-3(3x-12)=-13

OpenStudy (luigi0210):

Now solve: 2x-9x+36=-13 -7x=-39 x=39/7

OpenStudy (luigi0210):

*-7x=-49 x=7 Now plug x back into an equation and solve for y